# Solving PDEs in One Shot via Fourier Features with Exact Analytical Derivatives

### v1

@@ 3.4 Ablation Study
Table 4: Ablation study on Newton–Fast-LSQ components. Each row removes one component from the full method and reports the resulting $L^{2}$ error on two representative problems. “Diverged” indicates that Newton iteration failed to converge within 60 iterations. The full method uses all four components.
Configuration NL-Poisson 2D Steady Burgers 1D
@@
Without $1/\sqrt{N}$ normalization 4.2e-4 Diverged
Without Tikhonov ($\mu=0$) 3.8e-5 1.1e-6
Without warm-start (zero init) 8.3e-7 Diverged
Without line search (fixed $\alpha=1$) 9.4e-7 2.1e-5
Without continuation (Burgers only) Diverged
Table 4 isolates the contribution of each algorithmic component introduced in §2.4. The $1/\sqrt{N}$ normalization is the single most important factor: without it, the coefficients $\beta_{j}$ grow to $\mathcal{O}(10^{6})$$\mathcal{O}(10^{8})$ magnitudes, causing severe numerical cancellation in the residual evaluation and either degrading accuracy by four orders of magnitude (NL-Poisson) or causing outright divergence (Burgers). Tikhonov regularization ($\mu=10^{-10}$) contributes roughly three orders of magnitude of accuracy on NL-Poisson; even this tiny regularization stabilizes the Newton step direction in the near-nullspace of the Jacobian. The warm-start from the linear part provides a good initial guess, reducing both iteration count and the risk of convergence to spurious solutions; removing it increases the NL-Poisson error by an order of magnitude and causes Burgers to diverge because the nonlinear advection term dominates from the outset. The backtracking line search contributes about one order of magnitude of accuracy improvement by preventing overshooting in the later Newton iterations where the Jacobian is poorly conditioned. For Steady Burgers, continuation is indispensable: without gradually decreasing viscosity, Newton iteration at $\nu=0.1$ diverges regardless of the other components.
@@ 3.5 Nonlinear PDE Results: Regression Mode
Table 5: Nonlinear PDEs: function approximation (regression mode) with frozen features. Both Fast-LSQ and PIELM fit known exact solutions. PINNacle errors and times are from the best PINN variant (full PDE solve, not regression).
@@
Table 5 compares function approximation quality for nonlinear PDE solutions. Sinusoidal bases excel on smooth and oscillatory solutions: Klein-Gordon ($4.6\times 10^{-7}$), Navier-Stokes ($4.9\times 10^{-7}$), and especially Sine-Gordon ($3.4\times 10^{-4}$ vs. $5.5\times 10^{-2}$, a $160\times$ improvement over $\tanh$). The $\tanh$ basis shows comparable or slightly better performance only on the Burgers shock, which contains a sharp discontinuity better matched to sigmoid-type functions; even there, the gradient error favors $\tanh$ ($5.6\times 10^{-3}$ vs. $2.6\times 10^{-2}$) because the sharp transition in the shock layer is better captured by the saturating nonlinearity.
Where PINNacle comparisons are available, even our regression baseline (pure function fitting, not PDE solving) matches or exceeds the best PINN PDE solvers. On Burgers, $1.3\times 10^{-3}$ vs. $1.3\times 10^{-2}$ ($10\times$ better); on Gray-Scott, $8.3\times 10^{-6}$ vs. $8.0\times 10^{-2}$ (nearly $10{,}000\times$ better in under 0.1 s versus $\sim$612 s). These results underscore that the representational power of sinusoidal random features is itself a major advantage, independent of how the coefficients are optimized.
@@ 3.6 Gradient Accuracy
Table 6: Value and gradient relative $L^{2}$ errors for linear PDEs (solver mode).
Fast-LSQ (sin) PIELM (tanh)
Problem Value Gradient Value Gradient
Poisson 5D 4.8e-7 4.9e-6 4.7e-6 4.9e-5
Heat 5D 6.9e-4 3.9e-3 3.0e-3 1.1e-2
Wave 1D 1.3e-6 1.5e-6 1.8e-3 3.2e-3
Helmholtz 2D 1.9e-6 1.3e-6 7.4e-5 5.0e-5
Maxwell 2D 6.7e-7 1.0e-6 4.5e-5 1.1e-4
Accurate gradients are crucial for evaluating PDE residuals and computing downstream physical quantities such as fluxes and stresses. Table 6 reports both value and gradient errors for the linear PDEs. For Fast-LSQ, gradient errors are within one order of magnitude of value errors across all five problems, confirming high derivative quality. This is a direct consequence of (2): since $\nabla\phi_{j}$ involves the same random weights scaled by $\mathcal{O}(\sigma)$, the gradient inherits the approximation quality of the value fit. For PIELM, the value-to-gradient error ratio is comparable, but the absolute gradient errors are $\times100×1$$0\times$ worse due to the inferior value approximation.
@@ 3.7 Spectral Sensitivity Analysis
The bandwidth $\sigma$ controls the frequency content of the random Fourier features and must be matched to the solution’s spectral properties. Our grid search across $\sigma\in\{0.5,1,2,3,5,8,10,12,15\}$ reveals a clear pattern. Oscillatory problems (Wave, Helmholtz, Maxwell) prefer high $\sigma$ (between 5 and 15) to resolve the dominant frequency. Smooth problems (Poisson, Klein-Gordon) prefer moderate $\sigma$ (between 1 and 5) where the kernel length scale matches the solution’s smoothness. Discontinuous problems (Burgers shock) favor low-to-moderate $\sigma$ for sinusoidal bases, because high frequencies produce Gibbs ringing near the shock; $\tanh$ is less sensitive in this regime. Multiscale problems benefit from the multi-block architecture, where different blocks can target different scales. Full sensitivity plots for all problems are provided in Appendix C.
@@ 4 Related Work
PIELM (Dwivedi and Srinivasan, 2020) pioneered the frozen-feature linear-solve approach for linear PDEs using $\tanh$ activations. Like Fast-LSQ, PIELM solves a single least-squares system and is not iterative; the key distinction is that PIELM assembles the operator matrix via automatic differentiation because $\tanh$ derivatives lack a closed-form cyclic structure. As Proposition 2.2 shows, the symbolic expansion of $\frac{d^{n}}{dz^{n}}\tanh(z)$ grows as $\mathcal{O}(2^{n})$ terms, making analytic assembly impractical. In our experiments, PIELM is $10\times$$1000\times$ less accurate than Fast-LSQ at equal feature counts, demonstrating that the basis function choice has a profound effect on solution quality even when the solve method is identical. RF-PDE (Liao, 2024) handles both linear and nonlinear PDEs but uses iterative nonlinear least-squares optimization requiring 600–2000 epochs, despite the fact that the linear case admits a closed-form solution. Liao (2024) propose related random feature PDE solvers with different optimization strategies.
@@ 6 Conclusion
We presented Fast-LSQ, a method combining the one-shot linear solve of frozen-feature models with the exact cyclic derivative structure unique to sinusoidal bases. On linear PDEs, this yields $10^{-7}$ accuracy in under 0.1 s, orders of magnitude faster and more accurate than PINNs ($10^{-2}$$10^{-4}$ in 270–7500 s), RF-PDE ($10^{-4}$ with iterative optimization), and $\tanh$-based alternatives ($10^{-5}$$10^{-3}$ with identical solve protocol). The Newton–Raphson extension achieves $10^{-8}$ to $10^{-9}$ on nonlinear PDEs in under 9 s, sometimes outperforming regression oracles that use exact solutions. The ablation study shows that $1/\sqrt{N}$ normalization and Tikhonov regularization are essential for robust convergence, and continuation is indispensable for advection-dominated problems. Future directions include adaptive frequency selection, preconditioning for high-order PDEs, extension to vector-valued PDE systems, and a complete theoretical analysis of approximation, discretization, and conditioning bounds.
@@
Code and all experimental scripts will be made publicly available upon acceptance.
@@ Appendix A Detailed Comparison with RF-PDE
Appendix B Convergence Profiles
@@
Appendix C Spectral Sensitivity Plots

### v3

@@ 3.4 Ablation Study
Table 3: Ablation study. Each row removes one component; “Div.” = diverged in 60 iterations.
Configuration NL-Poisson Burgers
@@
No $1/\!\sqrt{N}$ norm 4.2e-4 Div.
No Tikhonov ($\mu{=}0$) 3.8e-5 1.1e-6
No warm-start 8.3e-7 Div.
No line search 9.4e-7 2.1e-5
No continuation Div.
Table 4: Value and gradient $L^{2}$ errors on linear PDEs (solver mode).
Fast-LSQ PIELM
Problem Val Grad Val Grad
Poisson 5D 4.8e-7 4.9e-6 4.7e-6 4.9e-5
Heat 5D 6.9e-4 3.9e-3 3.0e-3 1.1e-2
Wave 1D 1.3e-6 1.5e-6 1.8e-3 3.2e-3
Helmholtz 2D 1.9e-6 1.3e-6 7.4e-5 5.0e-5
Maxwell 2D 6.7e-7 1.0e-6 4.5e-5 1.1e-4
Table 4 isolates the contribution of each component from §2.4. The $1/\sqrt{N}$ normalization is critical: without it, coefficients grow to $\mathcal{O}(10^{6}--10^{8})$, causing either $4$ orders of magnitude accuracy loss or divergence. Tikhonov regularization, warm-start, and line search each contribute 1–3 orders of magnitude; continuation is indispensable for Burgers.
@@ 3.5 Nonlinear PDE Results: Regression Mode
Table 5: Nonlinear PDEs: regression mode (fitting known solutions). PINNacle: best PINN variant (full PDE solve, not regression).
@@
Table 5 compares function approximation quality. Sinusoidal features excel on smooth and oscillatory solutions: Sine-Gordon ($3.4\times 10^{-4}$ vs. $5.5\times 10^{-2}$, $160\times$ improvement), Klein-Gordon and Navier-Stokes at $\sim\!10^{-7}$. The $\tanh$ basis is comparable only on the Burgers shock, whose discontinuity suits sigmoid functions. Where PINNacle comparisons exist, our regression baseline alone (pure fitting, no PDE) exceeds the best PINN solvers: Gray-Scott $8.3\times 10^{-6}$ vs. $8.0\times 10^{-2}$ ($10{,}000\times$ better, 0.1 s vs. 612 s).
@@ 3.6 Gradient Accuracy
Table 4 reports value and gradient errors for the linear PDEs. Because the derivatives of a sinusoidal basis are just phase-shifted sinusoids scaled by $\mathcal{O}(\sigma)$, Fast-LSQ’s gradient errors remain tightly bounded: typically within one order of magnitude of its value errors across all five problems. In contrast, PIELM’s gradient errors are $\times100×1$$0\times$ worse than its value errors, largely due to $\tanh$’s inferior underlying value approximation and lack of the eigenfunction structure that keeps trigonometric derivatives well-behaved.
@@ 3.7 Spectral Sensitivity Analysis
The bandwidth $\sigma$ controls the frequency content of the random Fourier features and must be matched to the solution’s spectral properties. Our grid search across $\sigma\in\{0.5,1,2,3,5,8,10,12,15\}$ reveals a clear pattern. Oscillatory problems (Wave, Helmholtz, Maxwell) prefer high $\sigma$ (between 5 and 15) to resolve the dominant frequency. Smooth problems (Poisson, Klein-Gordon) prefer moderate $\sigma$ (between 1 and 5) where the kernel length scale matches the solution’s smoothness. Discontinuous problems (Burgers shock) favor low-to-moderate $\sigma$ for sinusoidal bases, because high frequencies produce Gibbs ringing near the shock; $\tanh$ is less sensitive in this regime. Multiscale problems benefit from the multi-block architecture, where different blocks can target different scales. Full sensitivity plots for all problems are provided in Appendix D.
3.8 Downstream Applications: Equation Discovery and Inverse Problems
The sub-second solve time and closed-form derivative access of Fast-LSQ unlock applications that are impractical with iterative PDE solvers. We briefly illustrate two: PDE discovery from data and inverse-problem parameter recovery.
PDE discovery via sparse regression.
SINDy (Brunton et al., 2016) discovers governing equations from data by building a dictionary of candidate terms and applying sparse regression. The standard approach computes derivatives via finite differences, which amplifies measurement noise: at noise level $\sigma_{\varepsilon}{=}0.01$ on $u$, the central-difference $u_{xx}$ has RMSE $\sim\!2500$ whereas the analytical derivatives from (2) achieve RMSE $\sim\!0.4$, a $\sim\!6000\times$ cleaner signal (Figure 2a). This dramatically extends the noise regime in which SINDy can operate. On 2000 observations of a damped oscillator ($u=e^{-x/2}\sin 2x$), LASSO with the analytical dictionary correctly identifies $u_{xx}=c_{1}u+c_{2}u_{x}$, recovering the stiffness coefficient to $0.2\%$ and the damping to $\sim\!12\%$ error, while driving the spurious $u^{2}$ term to zero (Figure 2; details in Appendix G.1).
Inverse heat-source localisation.
Because the forward mapping “source parameters $\to$ predicted temperature” passes through a single pre-factored linear solve $\boldsymbol{\beta}=(\mathbf{A}^{\top}\mathbf{A}+\mu I)^{-1}\mathbf{A}^{\top}\mathbf{f}$, gradients with respect to the source parameters are available analytically. We demonstrate this on a challenging multi-source problem: recovering four unknown anisotropic Gaussian heat sources (24 parameters total: position, intensity, and shape per source) from just 4 irregularly placed sensors, each recording a temperature time-series over 40 snapshots—480 observations in total accumulating as the space-time field $u(x,y,t)$ evolves. The Cholesky pre-factored forward solve is assembled once; each L-BFGS-B evaluation then costs only two triangular back-substitutions, making 150 iterations of 24-parameter optimization feasible in under a minute on CPU. Three of the four source positions are recovered to within 0.0010.0010.070.07 per coordinate; the fourth (Source 4, weakest source) has a larger $x$-error of $0.12$, reflecting the fundamental information limit of 4 sparse sensors for a 24-parameter problem (Figure 3, Table 6).
Sparse-sensor coil localisation.
A second inverse problem recovers the location of hidden current coils in a variable-permeability quadrupole magnet from 8 sparse magnetic-field measurements. The magnetic vector potential satisfies $\nabla\cdot(\nu(x,y)\,\nabla A_{z})=-J(x,y)$, a variable-coefficient PDE; expanding gives $\nu\,\Delta A_{z}+\nabla\nu\cdot\nabla A_{z}=-J$. Both $\Delta\phi_{j}$ and $\nabla\phi_{j}$ are evaluated analytically via (2)—no automatic differentiation is required even for this variable-coefficient case. The system matrix is assembled once and Cholesky-prefactored; each of the 40 Adam steps costs only two triangular back-substitutions. Starting from initial guess $(0.2,0.8)$, the true coil position $(0.45,0.45)$ is recovered to within $0.02$ in both coordinates (Figure 4).
Figure 2: PDE discovery with analytical derivatives ($\sigma_{\varepsilon}{=}0.01$). (a) Predicted $u_{xx}$ for each equation discovered at different LASSO sparsity levels; the OLS-refit on the selected support (red solid) closely tracks the truth (black). (b) LASSO penalty sweep; colour-matched dots mark the equation transitions.
Figure 3: Inverse heat-source localisation with 4 anisotropic Gaussian sources. Left: true temperature field at t=0.9t{=}0.9; white squares are the 4 irregularly placed sensors (each recording a temperature time-series over 40 snapshots); coloured stars are the 4 true source centres. Centre: recovered temperature field; coloured circles show the recovered centres (cf. stars). Right: L-BFGS-B optimisation trajectories for each source; triangles mark initial guesses, circles the converged estimates, faint stars the ground truth. All 4 source positions are recovered to within $0.06$ despite starting from incorrect locations.
Figure 4: Sparse-sensor coil localisation in a variable-$\mu$ quadrupole. Left: relative permeability $\mu_{r}(x,y)$; blue dashed circle marks the aperture; green diamonds are the 8 sparse sensors; white stars mark the four true hidden coil positions. Right: reconstructed $A_{z}$ (colour) and field lines (contours) after 40 Adam iterations; yellow stars show the recovered coil locations (agreement <0.02<0.02 in each coordinate).
@@ 4 Related Work
PIELM (Dwivedi and Srinivasan, 2020) pioneered the frozen-feature linear-solve approach for linear PDEs using $\tanh$ activations. Like Fast-LSQ, PIELM solves a single least-squares system and is not iterative; the key distinction is that PIELM requires handwritten, problem-specific symbolic calculus to assemble $\mathcal{L}[\tanh]$ for each new PDE operator, because $\tanh$ lacks a closed-form cyclic derivative structure. Fast-LSQ’s trigonometric features are operator-agnostic: one formula (2) applies to any linear differential operator (Proposition 2.2). Beyond this, the trigonometric formulation offers numerical advantages well-suited to PDEs: eigenfunction structure for constant-coefficient operators, orthogonality for spectral accuracy, and phase-shifted derivatives scaled by $\mathcal{O}(\sigma)$ that keep gradient errors tightly bounded. In our experiments, PIELM is $10\times$$1000\times$ less accurate than Fast-LSQ at equal feature counts, with gradient errors $\times100×1$$0\times$ worse; the basis function choice has a profound effect on solution quality even when the solve method is identical. RF-PDE (Liao, 2024) handles both linear and nonlinear PDEs but uses iterative nonlinear least-squares optimization requiring 600–2000 epochs, despite the fact that the linear case admits a closed-form solution. Liao (2024) propose related random feature PDE solvers with different optimization strategies.
RBF collocation.
Kansa’s method (Kansa, 1990) and its variants (Fasshauer, 2007) solve PDEs by collocation with radial basis functions (RBFs). Like Fast-LSQ, RBF Kansa is meshfree and one-shot (no iterative training), assembling a dense linear system by differentiating the RBF kernel analytically. The key distinction is the basis: RBFs such as multiquadric or Gaussian are governed by a shape parameter $c$ whose choice trades accuracy against conditioning, and achieving $10^{-4}$ errors typically requires careful tuning. Sinusoidal random features lack a shape-parameter sensitivity problem—bandwidth $\sigma$ is selected once by grid search—and achieve $10^{-8}$ errors where comparable MQ-RBF achieves $10^{-5}$ at equal DoF. Furthermore, RBF Kansa fails to solve stiff 1-D problems (Burgers $\nu{=}0.1$) and high-frequency Helmholtz (k=10k{=}10), while Fast-LSQ handles both without modification (Appendix G.4).
@@ 6 Conclusion
We presented Fast-LSQ, a method combining the one-shot linear solve of frozen-feature models with the exact cyclic derivative structure unique to sinusoidal bases. On linear PDEs, this yields $10^{-7}$ accuracy in under 0.1 s, orders of magnitude faster and more accurate than PINNs ($10^{-2}$$10^{-4}$ in 270–7500 s), RF-PDE ($10^{-4}$ with iterative optimization), and $\tanh$-based alternatives ($10^{-5}$$10^{-3}$ with identical solve protocol). The Newton–Raphson extension achieves $10^{-8}$ to $10^{-9}$ on nonlinear PDEs in under 9 s, sometimes outperforming regression oracles that use exact solutions. The ablation study shows that $1/\sqrt{N}$ normalization and Tikhonov regularization are essential for robust convergence, and continuation is indispensable for advection-dominated problems. The analytical framework enables several further extensions, validated in §3.8 and Appendix G. Learnable bandwidth (Appendices E and G.5): reparameterising $\mathbf{W}_{j}=\sigma\hat{\mathbf{W}}_{j}$ makes $\sigma$ differentiable through the exact inner solve; on Helmholtz 2D the loss drops five orders of magnitude in 80 AdamW steps. Inverse problems3.8): gradients flow through the pre-factored solve, enabling simultaneous recovery of 4 anisotropic heat sources (24 parameters) from just 4 irregularly placed sensors recording 240 observations, and sparse-sensor coil localisation in a variable-permeability quadrupole magnet (position error <0.02<0.02) in 40 Adam steps. Matrix caching (Appendix G.5): pre-computing $\mathbf{A}^{\dagger}$ yields a $362\times$ speedup for boundary-condition sweeps. PDE discovery3.8): the analytical derivative dictionary provides $\sim\!6000\times$ cleaner second derivatives than finite differences, extending the operational noise regime of SINDy-style sparse regression (Brunton et al., 2016).
Further directions.
Preconditioning for high-order PDEs, extension to vector-valued systems, domain decomposition for complex geometries, and a rigorous theoretical analysis of approximation and conditioning bounds remain important open problems.
@@
Full code including all examples and baselines is publicly available at https://github.com/sulcantonin/FastLSQ as a universal framework for Fast-LSQ with demos. The package is also available via pip install fastlsq. Code examples are provided in Appendix B.
@@ Appendix A Detailed Comparison with RF-PDE
Appendix B Code Examples
The Fast-LSQ framework provides minimal APIs for linear, nonlinear, and inverse problems. No analytical solution is required—the inverse example defines the PDE via Op, generates synthetic observations from the forward model, and recovers source parameters:
Listing 1: Linear, nonlinear, and inverse PDE solving (no exact solution needed).
1from fastlsq import solve_linear, solve_nonlinear, Op, solve_lstsq
2from fastlsq import SinusoidalBasis, sample_box, sample_boundary_box
3from fastlsq.problems.linear import PoissonND
4from fastlsq.problems.nonlinear import NLPoisson2D
5import torch
6import numpy as np
7from scipy.optimize import minimize
8
9# 1. Linear PDE
10result = solve_linear(PoissonND(), scale=5.0)
11
12# 2. Nonlinear PDE
13result = solve_nonlinear(NLPoisson2D(), max_iter=30)
14
15# 3. Inverse: recover source position (x_s,y_s) from sensor data
16# PDE: -Delta u = f, f = Gaussian at (x_s,y_s), u=0 on boundary
17pde_op = -Op.laplacian(d=2)
18basis = SinusoidalBasis.random(2, 800, sigma=5.0, normalize=True)
19x_pde = sample_box(3000, 2); x_bc = sample_boundary_box(400, 2)
20cache = basis.cache(x_pde)
21A_pde = pde_op.apply(basis, x_pde, cache=cache)
22A = torch.cat([A_pde, 100*basis.evaluate(x_bc)])
23x_sens = torch.tensor([[0.3,0.3],[0.7,0.7],[0.3,0.7],[0.7,0.3]])
24def fwd(xs, ys):
25 b = torch.exp(-((x_pde[:,0]-xs)**2+(x_pde[:,1]-ys)**2)/0.1).unsqueeze(1)
26 b = torch.cat([b, torch.zeros(400,1)])
27 beta = solve_lstsq(A, b)
28 return (basis.evaluate(x_sens) @ beta).detach().numpy().ravel()
29u_obs = fwd(0.4, 0.6) + 0.01*np.random.randn(4)
30def loss(p): return np.sum((fwd(p[0], p[1]) - u_obs)**2)
31minimize(loss, [0.5, 0.5], method="L-BFGS-B", bounds=[(0.1,0.9)]*2)
Appendix C Convergence Profiles
@@
Appendix D Spectral Sensitivity Plots